### v2

@@ IV.1.1 $D\bar{B}$ systems
Figure 4: $S$-wave potential of the $D\bar{B}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of $I(J^{P})=1(0^{+})$ $D\bar{B}$ system at $\Lambda=0.5~{}\rm{GeV}$, (b) the potential of $I(J^{P})=0(0^{+})$ $D\bar{B}$ system at $\Lambda=0.5~{}\rm{GeV}$, (c) the total potential of $D\bar{B}[1(0^{+})]$ system with different cutoffs. The green dashed, red dashed lines, and purple solid lines represent the contact, TPE, and full potential, respectively.
@@
It is worth noting that the effective potentials of the $I(J^{P})=1(0^{+})$ $D\bar{B}$ system have similar behaviors with the potentials of the $I(J^{P})=1(0^{+})$ $\bar{B}\bar{B}$ system in Ref. Wang:2018atz , which is a consequence of heavy flavor symmetry between the charm and bottom quarks. As shown in Fig. 4(c), the total potentials become more and more attractive as the cutoff parameter $\Lambda$ increases. The bound states emerge at $\Lambda=0.7~{}\rm{GeV}$ and $\Lambda=1.0~{}\rm{GeV}$, lying $0.8~{}\rm{MeV}$ and $3.6~{}\rm{MeV}$ below the $D\bar{B}$ mass threshold, respectively. However, we change the cutoff to $0.54~{}\rm{GeV}$, the very shallow bound state can emerge in the $D\bar{B}[0(0^{+})]$ channel.
@@ IV.1.2 $D\bar{B}^{*}$ systems
In this section, we investigate the interactions of the $D\bar{B}^{*}$ systems with $I(J^{P})=1(1^{+})$ and $I(J^{P})=0(1^{+})$. As shown in Fig. 2 before, only the contact diagram contributes to the amplitudes at LO. The one-loop corrections to the contact diagrams and TPE diagrams appear at NLO. Their results are shown in Figs. 5(a) and 5(b) we adopt the cutoff parameter $\Lambda$ with $0.5~{}\rm{GeV}$.
Comparing Fig. 5(a) with Fig. 4(a), we observe that the behavior of the $D\bar{B}^{*}[1(1^{+})]$ effective potential is similar with the $D\bar{B}[1(0^{+})]$ effective potential. For the $D\bar{B}^{*}[1(1^{+})]$ system, the repulsive contact interaction is much stronger than the attractive TPE interaction, which leads to a net repulsive potential. As a result, no bound state exists in this system.
In the case of the $D\bar{B}^{*}[0(1^{+})]$ system, the effective potentials are shown in Fig. 5(b). Compared to the 1($1(1^{+})$ system, the behaviors of the $0(1^{+})$ contact and TPE potential are both reversed, which results in an overall attractive interaction. However, this potential is weaker even than the one in the previous $D\bar{B}[0(0^{+})]$ system. As shown in Fig. 5(c), the total potential at $\Lambda=0.7~{}\rm{GeV}$ and at $\Lambda=0.5~{}\rm{GeV}$ are of similar strength. However, when $\Lambda=1.0~{}\rm{GeV}$, the effective potential is repulsive near the origin but becomes attractive at intermediate distances. This behavior appears because the TPE contribution at large momentum is more sensitive to cutoff $\Lambda$ variations, which leads to a repulsive potential at short distances. In our calculation, no bound state is found in the $D\bar{B}^{*}[0(1^{+})]$ system.
Figure 5: $S$-wave potential of the $D\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of $I(J^{P})=1(1^{+})$ $D\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (b) the potential of $I(J^{P})=0(1^{+})$ $D\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (c) the total potential of $D\bar{B}^{*}[0(1^{+})]$ system with different cutoffs. The green dashed and red dashed lines represent the contact and TPE contributions, respectively. The purple solid lines stand for the full potential.
Therefore we conclude that the interaction in the $D\bar{B}^{*}[0(1^{+})]$ system is weaker than that in the $D\bar{B}[0(0^{+})]$ system. This is consistent with the predictions of the one-boson-exchange potential model Li:2012ss , but contrasts with the conclusions drawn from lattice QCD calculations Alexandrou:2023cqg .
@@ IV.1.3 $D^{*}\bar{B}^{*}$ systems
In this section, we explore the possibility of bound states in the $D^{*}\bar{B}^{*}$ systems, which includes six different isospin-spin states: $1(0^{+})$, 0($0(0^{+})$, 1($1(1^{+})$, $0(1^{+})$, $1(2^{+})$, and $0(2^{+})$. At LO, both contact and OPE contribute to the effective potential. At NLO, the one-loop corrections to the contact and OPE diagrams shown in Fig. 3 as well as TPE diagrams all contribute.
First, we focus on the $J^{P}=0^{+}$ $D^{*}\bar{B}^{*}$ state. The effective potentials are shown in Figs. 6(a) and 6(b), we adopt the cutoff parameter $\Lambda$ with $0.5~{}\rm{GeV}$. In Fig. 6(a), we observe that both the contact and OPE interactions are repulsive, while the TPE interaction is attractive. However, the total potential is repulsive, meaning no bound state can exist in the $1(0^{+})$ $D^{*}\bar{B}^{*}$ system. In contrast, for the 0($0(0^{+})$ system in Fig. 6(b), the contact and OPE interactions provide an attractive force, while the TPE interaction is repulsive. The repulsive TPE contribution partially counteracts the attractive contact interaction, but the OPE interaction remains strong enough to form a bound state. The resulting binding energy is $\Delta E=6.0~{}\rm{MeV}$. As shown in Fig. 6, the OPE terms can contribute to the total interaction of $D^{*}\bar{B}^{*}$ systems, it is a departure from the $D\bar{B}^{(*)}$ systems.
@@
Next, we analyze the interactions in the $D^{*}\bar{B}^{*}[1^{+}]$ system. For the 1($1(1^{+})$ state, as seen in Fig. 7(a), the OPE contribution is largely offset by the TPE contribution, with the repulsive contact potential remaining. It results in a fully repulsive total potential, meaning no bound state exists in the $D^{*}\bar{B}^{*}[1(1^{+})]$ system. For the $D^{*}\bar{B}^{*}[0(1^{+})]$ system, the effective potentials, shown in Fig. 7(b), exhibit similar behaviors to those of the $D^{*}\bar{B}^{*}[0(0^{+})]$ system. Although the attraction is not as strong as in the $D^{*}\bar{B}^{*}[0(0^{+})]$, it is still sufficient to form a shallow bound state. By solving the Schrödinger equation, we obtain the $D^{*}\bar{B}^{*}[0(1^{+})]$ binding energy $\Delta E=0.6~{}\rm{MeV}$ at $\Lambda=0.5~{}\rm{GeV}$.
Figure 6: $S$-wave potential of $J=0$ $D^{*}\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of $I(J^{P})=1(0^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (b) the potential of $I(J^{P})=0(0^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (c) the total potential of $D^{*}\bar{B}^{*}[0(0^{+})]$ system with different cutoffs. The green dashed, blue dashed, and red dashed lines represent the contact, OPE, and TPE contributions, respectively. The purple solid lines stand for the full potential.
Figure 7: $S$-wave potential of $J=1$ $D^{*}\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of $I(J^{P})=1(1^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (b) the potential of $I(J^{P})=0(1^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (c) the total potential of $D^{*}\bar{B}^{*}[0(1^{+})]$ system on the different cutoffs. The green dashed, blue dashed, and red dashed lines represent the contact, OPE, and TPE contributions, respectively. The purple solid lines stand for the full potential.
Figure 8: $S$-wave potential of $J=2$ $D^{*}\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of $I(J^{P})=1(2^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (b) the potential of $I(J^{P})=0(2^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~{}\rm{GeV}$, (c) the total potential of $D^{*}\bar{B}^{*}[0(2^{+})]$ system with the different cutoffs. The green dashed, blue dashed, and red dashed lines represent the contact, OPE, and TPE contributions, respectively. The purple solid lines stand for the full potential.
@@
Besides, in the $D^{*}\bar{B}^{*}[0(0^{+})]$ and $D^{*}\bar{B}^{*}[0(1^{+})]$ channels, two shallow bound states are found with binding energies of $6.0~{}\rm{MeV}$ and $0.6~{}\rm{MeV}$, respectively. Therefore, we suggest that future theoretical and experimental studies should focus on these two systems to find new heavy tetraquark states. Compared to the total potentials of the five $I=0$ channels, we can also conclude that the interaction in $D^{*}\bar{B}^{*}[0(0^{+})]$ channel is more attractive than the other four channels. The mass and the binding energy of each $D^{(*)}\bar{B}^{(*)}[I=0]$ system under different cutoffs are listed in Table 9.
@@
$0.5$cdots $\cdots$ $\cdots$ $\cdots$ $7235.7$ $6.0$ $7241.1$ $0.6$cdots $\cdots$
$0.7$ $7143.6$ $0.8$cdots $\cdots$ $7230.6$ $11.1$ $7240.0$ $1.7$cdots $\cdots$
$1.0$ $7140.8$ $3.6$cdots $\cdots$ $7205.8$ $35.9$ $7232.3$ $9.4$ $7237.4$ $4.3$
@@ IV.2 Coupled-channel effects
We have considered interactions of the $D^{(*)}\bar{B}^{(*)}$ systems in the single-channel case. In this section, we will investigate how much their coupled-channel effects affect the above results. We categorize $I=0$ coupled-channel $D^{(*)}\bar{B}^{(*)}$ systems into two groups according to their $0(J^{P})$ quantum number: $D\bar{B}(^{1}S_{0})$, $D^{*}\bar{B}^{*}(^{1}S_{0})$}$[0(0^{+})]$ and $D\bar{B}^{*}(^{3}S_{1})$, $D^{*}\bar{B}(^{3}S_{1})$, $D^{*}\bar{B}^{*}(^{3}S_{1})$}$[0(1^{+})]$.
We first consider the $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ system composed of $D\bar{B}(^{1}S_{0})$, $D^{*}\bar{B}^{*}(^{1}S_{0})$}. In this case, the effective potentials of the inelastic scattering process $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{*}\bar{B}^{*}\to D\bar{B}$ are needed. We depict the $D\bar{B}\to D^{*}\bar{B}^{*}$ diagrams in Fig. 12. At LO, the contact and OPE diagrams contribute to the scattering amplitudes. At NLO, there are one-loop corrections to these LO diagrams and also TPE diagrams. The calculated scattering amplitudes for these diagrams are listed in Appendix A.1. The $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ coupled-channel potential in the coordinate space is
@@
where the inelastic scattering $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{*}\bar{B}^{*}\to D\bar{B}$ appear as non-diagonal elements in the $2\times 2$ potential matrix. Then we solve the coupled-channel Schrödinger equation and obtain the numerical results including the binding energy $E$, and the probability of the individual channel $P_{i}$, which are shown in Table 10.
Here we consider the variation on the cutoff $\Lambda$ just as in the single-channel case. Compared with the single-channel $D\bar{B}[0(0^{+}]$ results in Table 9, we find coupled channels somewhat strengthen the binding energy: from 0.8 GeV in the single channel to 1.1 GeV in the coupled channels at $\Lambda=0.7$ GeV, for example. Also, because of the large mass gap (about $190~{}{\rm{MeV}}$) between the $D\bar{B}$ and $D^{*}\bar{B}^{*}$ thresholds, the contribution of $D^{*}\bar{B}^{*}(^{1}S_{0})$ is $0.1\%$ at $\Lambda=0.5~{}\rm{GeV}$. If we raise the cutoff to $0.7~{}\rm{GeV}$ and $1.0~{}\rm{GeV}$, the probability of $D^{*}\bar{B}^{*}(^{1}S_{0})$ increases to $0.3\%$ and $0.9\%$, respectively.
@@
We still take the value of the cutoff parameter at $0.5~{}\rm{GeV}$, $0.7~{}\rm{GeV}$, and $1.0~{}\rm{GeV}$ to evaluate binding energy and probability $P_{i}$, respectively. The numerical results are listed in Table 11. We find no bound state in the $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ system which is consistent with the single-channel $D\bar{B}^{*}[0(1^{+})]$ calculation. The dominant channel is $D\bar{B}^{*}(^{3}S_{1})$, with a probability of $99.9\%-99.5\%$.
@@
Table 10: The numerical results for $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ systems in $S$-wave. The cutoff parameter $\Lambda$, and binding energy $E$ relative to the threshold of $D\bar{B}$ are in units of $\rm{GeV}$, and $\rm{MeV}$ respectively. $\cdots$” means that there is no bound solution.
@@
Table 11: The numerical results for $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ systems in $S$-wave. The cutoff parameter $\Lambda$, and binding energy $E$ relative to the threshold of $D\bar{B}^{*}$ are in units of $\rm{GeV}$, and $\rm{MeV}$ respectively. $\cdots$” means that there is no bound solution.

### v3

@@ IV.1.1 $D\bar{B}$ systems
Figure 4: $S$-wave potential of the $D\bar{B}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of the $I(J^{P})=1(0^{+})$ $D\bar{B}$ system at $\Lambda=0.5~\rm{GeV}$, (b) the potential of the $I(J^{P})=0(0^{+})$ $D\bar{B}$ system at $\Lambda=0.5~\rm{GeV}$, and (c) the total potential of the $D\bar{B}[1(0^{+})]$ system with different cutoffs. The green dashed, red dashed, and purple solid lines represent the contact, TPE, and full potential, respectively.
@@
It is worth noting that the effective potentials of the $I(J^{P})=1(0^{+})$ $D\bar{B}$ system have similar behaviors with the potentials of the $I(J^{P})=1(0^{+})$ $\bar{B}\bar{B}$ system in Ref. Wang:2018atz , which is a consequence of heavy flavor symmetry between the charm and bottom quarks. As shown in Fig. 4(c), the total potentials become more and more attractive as the cutoff parameter $\Lambda$ increases. The bound states emerge at $\Lambda=0.7~\rm{GeV}$ and $\Lambda=1.0~\rm{GeV}$, lying $0.8~\rm{MeV}$ and $3.6~\rm{MeV}$ below the $D\bar{B}$ mass threshold, respectively. However, if we change the cutoff to $0.54~\rm{GeV}$, the very shallow bound state can emerge in the $D\bar{B}[0(0^{+})]$ channel.
@@ IV.1.2 $D\bar{B}^{*}$ systems
In this section, we investigate the interactions of the $D\bar{B}^{*}$ systems with $I(J^{P})=1(1^{+})$ and $I(J^{P})=0(1^{+})$. As shown in Fig. 2 before, only the contact diagram contributes to the amplitudes at LO. The one-loop corrections to the contact diagrams and TPE diagrams appear at NLO. Their results are shown in Figs. 5(a) and 5(b), where we adopt the cutoff parameter $\Lambda$ with $0.5~\rm{GeV}$.
Comparing Fig. 5(a) with Fig. 4(a), we observe that the behavior of the $D\bar{B}^{*}[1(1^{+})]$ effective potential is similar to the $D\bar{B}[1(0^{+})]$ effective potential. For the $D\bar{B}^{*}[1(1^{+})]$ system, the repulsive contact interaction is much stronger than the attractive TPE interaction, which leads to a net repulsive potential. As a result, no bound state exists in this system.
In the case of the $D\bar{B}^{*}[0(1^{+})]$ system, the effective potentials are shown in Fig. 5(b). Compared to the $1(1^{+})$ system, the behaviors of the $0(1^{+})$ contact and TPE potential are both reversed, which results in an overall attractive interaction. However, this potential is weaker even than the one in the previous $D\bar{B}[0(0^{+})]$ system. As shown in Fig. 5(c), the total potentials at $\Lambda=0.7~\rm{GeV}$ and at $\Lambda=0.5~\rm{GeV}$ are of similar strength. However, when $\Lambda=1.0~\rm{GeV}$, the effective potential is repulsive near the origin but becomes attractive at intermediate distances. This behavior appears because the TPE contribution at large momentum is more sensitive to cutoff $\Lambda$ variations, which leads to a repulsive potential at short distances. In our calculation, no bound state is found in the $D\bar{B}^{*}[0(1^{+})]$ system.
Figure 5: $S$-wave potential of the $D\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of the $I(J^{P})=1(1^{+})$ $D\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, (b) the potential of the $I(J^{P})=0(1^{+})$ $D\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, and (c) the total potential of the $D\bar{B}^{*}[0(1^{+})]$ system with different cutoffs. The green dashed and red dashed lines represent the contact and TPE contributions, respectively. The purple solid lines stand for the full potential.
Therefore, we conclude that the interaction in the $D\bar{B}^{*}[0(1^{+})]$ system is weaker than that in the $D\bar{B}[0(0^{+})]$ system. This is consistent with the predictions of the one-boson-exchange potential model Li:2012ss , but contrasts with the conclusions drawn from lattice QCD calculations Alexandrou:2023cqg .
@@ IV.1.3 $D^{*}\bar{B}^{*}$ systems
In this section, we explore the possibility of bound states in the $D^{*}\bar{B}^{*}$ systems, which includes six different isospin-spin states: $1(0^{+})$, $0(0^{+})$, $1(1^{+})$, $0(1^{+})$, $1(2^{+})$, and $0(2^{+})$. At LO, both contact and OPE contribute to the effective potential. At NLO, the one-loop corrections to the contact and OPE diagrams shown in Fig. 3 as well as TPE diagrams all contribute.
First, we focus on the $J^{P}=0^{+}$ $D^{*}\bar{B}^{*}$ state. The effective potentials are shown in Figs. 6(a) and 6(b), where we adopt the cutoff parameter $\Lambda$ with $0.5~\rm{GeV}$. In Fig. 6(a), we observe that both the contact and OPE interactions are repulsive, while the TPE interaction is attractive. However, the total potential is repulsive, meaning no bound state can exist in the $1(0^{+})$ $D^{*}\bar{B}^{*}$ system. In contrast, for the $0(0^{+})$ system in Fig. 6(b), the contact and OPE interactions provide an attractive force, while the TPE interaction is repulsive. The repulsive TPE contribution partially counteracts the attractive contact interaction, but the OPE interaction remains strong enough to form a bound state. The resulting binding energy is $\Delta E=6.0~\rm{MeV}$. As shown in Fig. 6, the OPE terms can contribute to the total interaction of $D^{*}\bar{B}^{*}$ systems, which is a departure from the $D\bar{B}^{(*)}$ systems.
@@
Next, we analyze the interactions in the $D^{*}\bar{B}^{*}[1^{+}]$ system. For the $1(1^{+})$ state, as seen in Fig. 7(a), the OPE contribution is largely offset by the TPE contribution, with the repulsive contact potential remaining. It results in a fully repulsive total potential, meaning no bound state exists in the $D^{*}\bar{B}^{*}[1(1^{+})]$ system. For the $D^{*}\bar{B}^{*}[0(1^{+})]$ system, the effective potentials, shown in Fig. 7(b), exhibit similar behaviors to those of the $D^{*}\bar{B}^{*}[0(0^{+})]$ system. Although the attraction is not as strong as in the $D^{*}\bar{B}^{*}[0(0^{+})]$ system, it is still sufficient to form a shallow bound state. By solving the Schrödinger equation, we obtain the $D^{*}\bar{B}^{*}[0(1^{+})]$ binding energy $\Delta E=0.6~\rm{MeV}$ at $\Lambda=0.5~\rm{GeV}$.
Figure 6: $S$-wave potential of $J=0$ $D^{*}\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of the $I(J^{P})=1(0^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, (b) the potential of the $I(J^{P})=0(0^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, and (c) the total potential of the $D^{*}\bar{B}^{*}[0(0^{+})]$ system with different cutoffs. The green dashed, blue dashed, and red dashed lines represent the contact, OPE, and TPE contributions, respectively. The purple solid lines stand for the full potential.
Figure 7: $S$-wave potential of $J=1$ $D^{*}\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of the $I(J^{P})=1(1^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, (b) the potential of the $I(J^{P})=0(1^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, and (c) the total potential of the $D^{*}\bar{B}^{*}[0(1^{+})]$ system on the different cutoffs. The green dashed, blue dashed, and red dashed lines represent the contact, OPE, and TPE contributions, respectively. The purple solid lines stand for the full potential.
Figure 8: $S$-wave potential of $J=2$ $D^{*}\bar{B}^{*}$ systems and the dependence of the total potential on the cutoff parameter, namely, (a) the potential of the $I(J^{P})=1(2^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, (b) the potential of the $I(J^{P})=0(2^{+})$ $D^{*}\bar{B}^{*}$ system at $\Lambda=0.5~\rm{GeV}$, and (c) the total potential of the $D^{*}\bar{B}^{*}[0(2^{+})]$ system with the different cutoffs. The green dashed, blue dashed, and red dashed lines represent the contact, OPE, and TPE contributions, respectively. The purple solid lines stand for the full potential.
@@
Besides, in the $D^{*}\bar{B}^{*}[0(0^{+})]$ and $D^{*}\bar{B}^{*}[0(1^{+})]$ channels, two shallow bound states are found with binding energies of $6.0~\rm{MeV}$ and $0.6~\rm{MeV}$, respectively. Therefore, we suggest that future theoretical and experimental studies should focus on these two systems to find new heavy tetraquark states. Compared to the total potentials of the five $I=0$ channels, we can also conclude that the interaction in the $D^{*}\bar{B}^{*}[0(0^{+})]$ channel is more attractive than the other four channels. The mass and the binding energy of each $D^{(*)}\bar{B}^{(*)}[I=0]$ system under different cutoffs are listed in Table 9.
@@
$0.5$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $7235.7$ $6.0$ $7241.1$ $0.6$ $\cdots$ $\cdots$
$0.7$ $7143.6$ $0.8$ $\cdots$ $\cdots$ $7230.6$ $11.1$ $7240.0$ $1.7$ $\cdots$ $\cdots$
$1.0$ $7140.8$ $3.6$ $\cdots$ $\cdots$ $7205.8$ $35.9$ $7232.3$ $9.4$ $7237.4$ $4.3$
@@ IV.2 Coupled-channel effects
We have considered interactions of the $D^{(*)}\bar{B}^{(*)}$ systems in the single-channel case. In this section, we will investigate how much their coupled-channel effects affect the above results. We categorize the $I=0$ coupled-channel $D^{(*)}\bar{B}^{(*)}$ systems into two groups according to their $0(J^{P})$ quantum numbers: $D\bar{B}(^{1}S_{0})$, $D^{*}\bar{B}^{*}(^{1}S_{0})$$0(0^{+})]$ and $D\bar{B}^{*}(^{3}S_{1})$, $D^{*}\bar{B}(^{3}S_{1})$, $D^{*}\bar{B}^{*}(^{3}S_{1})$$0(1^{+})]$.
We first consider the $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ system composed of $D\bar{B}(^{1}S_{0})$, $D^{*}\bar{B}^{*}(^{1}S_{0})$}. In this case, the effective potentials of the inelastic scattering process $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{*}\bar{B}^{*}\to D\bar{B}$ are needed. We depict the $D\bar{B}\to D^{*}\bar{B}^{*}$ diagrams in Fig. 12. At LO, the contact and OPE diagrams contribute to the scattering amplitudes. At NLO, there are one-loop corrections to these LO diagrams and TPE diagrams. The calculated scattering amplitudes for these diagrams are listed in Appendix A.1. The $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ coupled-channel potential in the coordinate space is
@@
where the inelastic scattering $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{*}\bar{B}^{*}\to D\bar{B}$ appear as nondiagonal elements in the $2\times 2$ potential matrix. Then we solve the coupled-channel Schrödinger equation and obtain the numerical results, including the binding energy $E$, and the probability of the individual channel $P_{i}$, which are shown in Table 10.
Here, we consider the variation on the cutoff $\Lambda$ just as in the single-channel case. Compared with the single-channel $D\bar{B}[0(0^{+}]$ results in Table 9, we find coupled channels somewhat strengthen the binding energy: from 0.8 GeV in the single channel to 1.1 GeV in the coupled channels at $\Lambda=0.7$ GeV, for example. Also, because of the large mass gap (about $190~{\rm{MeV}}$) between the $D\bar{B}$ and $D^{*}\bar{B}^{*}$ thresholds, the contribution of $D^{*}\bar{B}^{*}(^{1}S_{0})$ is $0.1\%$ at $\Lambda=0.5~\rm{GeV}$. If we raise the cutoff to $0.7~\rm{GeV}$ and $1.0~\rm{GeV}$, the probability of $D^{*}\bar{B}^{*}(^{1}S_{0})$ increases to $0.3\%$ and $0.9\%$, respectively.
@@
We still take the values of the cutoff parameter at $0.5~\rm{GeV}$, $0.7~\rm{GeV}$, and $1.0~\rm{GeV}$ to evaluate binding energy and probability $P_{i}$, respectively. The numerical results are listed in Table 11. We find no bound state in the $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ system, which is consistent with the single-channel $D\bar{B}^{*}[0(1^{+})]$ calculation. The dominant channel is $D\bar{B}^{*}(^{3}S_{1})$, with a probability of $99.9\%-99.5\%$.
@@
Table 10: The numerical results for $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ systems in the $S$-wave. The cutoff parameter $\Lambda$ and binding energy $E$ relative to the threshold of $D\bar{B}$ are in units of $\rm{GeV}$ and $\rm{MeV}$, respectively. $\cdots$” means that there is no bound solution.
@@
Table 11: The numerical results for $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ systems in the $S$-wave. The cutoff parameter $\Lambda$ and binding energy $E$ relative to the threshold of $D\bar{B}^{*}$ are in units of $\rm{GeV}$ and $\rm{MeV}$, respectively. $\cdots$” means that there is no bound solution.